The first stock pays a current dividend of $2 which is growing at 3% per year. The cost of equity is 9% and the WACC is 7%. Based on this info what should you do if the stock is trading at $29.50 a share?

a. Buy it as the intrinsic value of the stock is just over $34

b. Short or sell the stock as the intrinsic value of the stock is under $26 a share

c. Buy it as the intrinsic value is $33.33

d. Buy is as the intrinsic value is just over $51
Answer #1

Value of Stock =P0=D1/(R-g)

D1=Next year's dividend

R=Required return on equity =9%=0.09

g=Constant growth rate =3%=0.03

D0=Current dividend=$2

D1=Next years dividend =$2*(1+0.03)=$2.06

Fair Market Value of Stock=P0=2.06/(0.09-0.03)=$34.33

The Stock Price of $29.50 is less than the fair market value of $34.33.

Hence, it should be bought.

ANSWER:

a. Buy it as the intrinsic value of the stock is just over $34